OpenStudy (anonymous):

Simplify the following expression. 4a-2b

OpenStudy (mathstudent55):

It is already simplified since the terms are unlike and can't be combined. It can be factored, though, into 2(2a - b)

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

can you also help me out with this one Evaluate the following expression. (81)0

OpenStudy (mathstudent55):

81 times zero?

OpenStudy (mathstudent55):

If so, remember that any number times zero equals zero, so (81)0 = 0

OpenStudy (anonymous):

sorry (8^1)^0

OpenStudy (mathstudent55):

Ok. 8 to the 1 power, then raised to the zero power?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

You can do this two ways: 1. First raise 8 to the 1 power. What is 8^1?

OpenStudy (anonymous):

so does it just equal 8?

OpenStudy (mathstudent55):

Yes. Any number raised to the 1 power is just the number. So (8^1)^0 = 8^0 Now what is a number raised to the zero power?

OpenStudy (anonymous):

1?

OpenStudy (mathstudent55):

Correct. The only exception is 0 because 0^0 is undefined, but any number, other than zero, raised to the zero power equals 1. So, we can continue: (8^1)^0 = 8^0 = 1

OpenStudy (mathstudent55):

That's the first way of solving this.

OpenStudy (mathstudent55):

The second way is using a rule of powers.

OpenStudy (mathstudent55):

When you raise an power to a power, you multiply the powers.

OpenStudy (mathstudent55):

The rule is (a^m)^n = a^(mn) For example, (2^3)^5 = 2^15 (multiply exponents 3 and 5 together)

OpenStudy (mathstudent55):

We can solve your problem with this rule: (8^1)^0 = 8^0 (since exponent 1 times exponent 0 = 0) Now we deal with 8^0 as a non-zero number raised tot eh zero poer: 8^0 = 1

OpenStudy (anonymous):

thanks so much! making it sound easier thank you!

OpenStudy (mathstudent55):

wlcm